Circular No. 6012

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

PSR J2043+27 S. E. Thorsett, P. S. Ray, S. R. Kulkarni, and T. A. Prince, California

Institute of Technology, report: “We have discovered a 96.13-ms pulsar in data taken on May 23 UT with the 305-m Arecibo telescope. The position is α = 20^{h}43^{m}.7, δ = +27^{o}40′ (equinox 2000.0; uncertainty of 5′ in radius). The flux density at 430 MHz is roughly 25 mJy. This could either be a young pulsar or a mildly recycled pulsar. We note that the pulsar is within a degree of the boundary of the southern blowout region of the Cygnus Loop supernova remnant, and the dispersion measure of 21 ± 1 pc cm^{−3} implies a distance of ∼ 1 kpc, consistent with the distance to the remnant. If the pulsar were born at the remnant center 40 kyr ago, its implied transverse velocity would be ∼ 1500 km/s.”

PULSAR IN G180.0–1.7 S. B. Anderson, B. J. Cadwell, A. Wolszczan, Penn State University;

and R. S. Foster, Naval Research Laboratory, report: “We have detected a 147-ms radio pulsar located near the center of the well-known supernova remnant G180.0–1.7 (S147). The new object was discovered on Apr. 3 during an ongoing pulsar survey with the 305-m Arecibo radio telescope at α = 5^{h}35^{m}.3, δ = +28^{o}11′ (equinox 1950.0); FWHM of the telescope at 430 MHz is 10′. Preliminary values of Ṗ = 4 × 10^{−15} and dispersion measure 40 pc cm^{−3} give the pulsar a spin-down age of 6 × 10^{5} yr and an estimated distance of 1.5 kpc, in agreement with the corresponding values of these parameters for S147. Together with the positional coincidence, this makes an association of the pulsar with S147 almost certain. Further timing and astrometric observations of the new object at Arecibo and the Very Large Array are in progress.”

NOVA SAGITTARII 1994 No. 2 L. V. Morrison and R. W. Argyle, Royal Greenwich Observatory, report

the following V magnitudes (cf. IAUC 6011; ± 0.05 mag): June 12.10 UT, 9.95; 13.10, 9.99; 14.09, 9.87; 15.09, 9.19; 16.09, 8.71; 17.08, 8.42; 18.08, 8.74; 20.08, 9.24; 22.07, 9.76; 27.06, 10.24.

1994 July 5 Daniel W. E. Green